### v3

# Cooperative Coevolution for Non-Separable Large-Scale Black-Box Optimization: Convergence Analyses and Distributed Accelerations

Abstract
Given the ubiquity of non-separable optimization problems in real worlds, in this paper we analyze and extend the large-scale version of the well-known cooperative coevolution (CC), a divide-and-conquer black-box optimization framework, on non-separable functions. First, we reveal empirical reasons of when decomposition-based methods are preferred or not in practice on some non-separable large-scale problems, which have not been clearly pointed out in many previous CC papers. Then, we formalize CC to a continuous-game model via simplification, but without losing its essential property. Different from previous evolutionary game theory for CC, our new model provides a much simpler but useful viewpoint to analyze its convergence, since only the pure Nash equilibrium concept is needed and more general fitness landscapes can be explicitly considered. Based on convergence analyses, we propose a hierarchical decomposition strategy for better generalization, as for any decomposition, there is a risk of getting trapped into a suboptimal Nash equilibrium. Finally, we use powerful distributed computing to accelerate it under the recent multi-level learning framework, which combines the fine-tuning ability from decomposition with the invariance property of CMA-ES. Experiments on a set of high-dimensional test functions validate both its search performance and scalability (w.r.t. CPU cores) on a clustering computing platform with 400 CPU cores.
keywords:
Black-box optimization , convergence analysis , cooperative coevolution , distributed algorithm , large-scale optimization
journal: arXiv
[…]
1 Introduction
Recent advances in artificial intelligence (particularly deep models [1, 2, 3]) and big data have generated a growing number of large-scale optimization (LSO) problems, including challenging high-dimensional black-box optimization (BBO) instances [4, 5]. An OpenAI team [6] used a highly-parallelized version of evolution strategies (ES) to optimize millions of weights of deep neural networks [7] for direct policy search in reinforcement learning (RL). In the recent Science paper, Fan et al. [8] hybridized genetic algorithm (GA) and simulated annealing (SA) to solve a big-data-based paleobiology model on a high-performing supercomputer called Tianhe. A DeepMind team proposed a population-based training method for hyper-parameter optimization of generative adversarial networks [9], multi-agent RL [10], and a deep vision model for self-driving cars [11].
Although evolutionary algorithms (EAs [12, 13, 14]) are one very popular algorithm family for BBO, nearly all standard versions of EAs have to be significantly improved in the LSO context, since they suffer easily from the notorious curse of dimensionality (from their random sampling operations). For a survey of EAs for LSO, see e.g., [15, 16, 17] and references wherein. For efficient search in large-scale space, exploiting the possibly useful problem structure is a critical step to accelerate convergence progress, as shown in Nesterov’s outstanding optimization book [18]. In this paper, we focus on mainly two of the representative EAs both with some successful LSO applications: 1) cooperative co-evolution (CC [19, 20, 21, 22, 23, 24]) based on the (co-adapted) modularity assumption; and 2) covariance matrix adaptation evolution strategies (CMA-ES [25, 26]), which is regarded widely as the state-of-the-art for BBO in a recent Nature review for EAs [13]. For modern CC, CMA-ES is often chosen as the basic suboptimizer for all subproblems, because of its well-studied theoretical properties (invariance against affine transformation [27], unbiases under neutral selection [28], maximum entropy (diversity) principle [29], learning of natural gradient [30, 31]) and generalizable search abilities in particular on non-separable and ill-conditioned optimization problems.
In 2014, a modern CC variant with the so-called differential grouping (DG) technique [32] obtained promising results on a class of partially additively separable (PAS) [33] benchmark functions for LSO (mainly from IEEE-CEC competitions [34, 35, 36]). Its well-established theoretical foundation to detect variable interactions has now sparked many following-up works (see Section 2.A). However, as previously argued in a Nature review [37], “most instances of most problems are not readily ‘linearly’ decomposable into building blocks”. Indeed, a plenty of real-world problems have complex (nonlinear) objective functions, where typically there have explicit or implicit interactions between any two variables for both gradient-based optimization [38] and BBO. Take e.g., [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] as examples, to name a few [^3] Here we provide an online website (https://tinyurl.com/3d4nnneb) which covers many real-world applications (nearly all of them are non-separable). What kinds of real-world applications are PAS is still an open question.. Even linear regression, perhaps the simplest data model, has a fully non-separable loss function form [^4] Only when the involved data has a block matrix structure, the resulting objective function is PAS. To our knowledge, however, little of real-world applications exhibit this (see https://archive.ics.uci.edu/ml/datasets.php for $>$ 150 data sets). In general, its objective function is fully non-separable (here we do not consider its complex dual form, which is clearly out of scope). [50] for least-squares estimation. In fact, in the original Ph.D. dissertation regarding CC [20], all of three real-world problems considered (i.e., string cover, rule learning, and neuroevolution) are non-separable, though CC was first benchmarked on both separable and non-separable artificially-constructed functions. Currently its state-of-the-art real-world applications come mainly from neuroevolution for RL (refer to Miikkulainen’s or Schmidhuber’s lab [51, 52, 53, 54, 55, 56, 57, 58]). Obviously, all the loss functions used by them are non-separable, caused by the nonlinearity of neural network itself as well as simulation model.
Given the ubiquity of non-separable problems in practice, in this paper we focus on CC for non-separable large-scale BBO mainly from three different yet related viewpoints (i.e., problem structure, convergence analyses, and distributed acceleration).
Problem Structure: It is natural to deduce that not all non-separable problems can be handled efficiently by CC. For example, on many ill-conditioned non-separable landscapes, both CC and its gradient-based counterpart (i.e., coordinate descent, CD [59]) are typically worse than second-order-type optimizers (e.g., LM-CMA [60] and L-BFGS [61]). On the contrary, on some other non-separable landscapes (e.g., with relatively sparse variable interactions [62]), some modern CC could obtain very competitive (sometimes even state-of-the-art) results. Therefore, a key theoretical question arises: On what kinds of problem types from (non-separable) real-world problems CC is preferred over others? As argued in 1989 by Conn et al. [63], PAS functions “are clearly a very restricted case of partially separable (PS) functions [^5] We notice that there are several CC papers do not distinguish PAS and PS clearly, which may cause confusions. In the mathematical optimization (MO) community, PS includes not only PAS, but also non-separable functions with a sparse linkage structure. In effect, the focus of MO is the latter rather than the former, as the former is seen as “a very restricted case”, which is not emphasized by some CC papers.. Here we exclude PAS because of its hard-to-satisfy assumption in practice.
To the best of our knowledge, there have no theoretical work to satisfactorily solve the above theoretical challenge. In this paper, we first answer a related but much simpler practical question: When may end-users prefer to use divide-and conquer methods in practice? Some common answers are presented in the following:
  • 1.
    There often exists a “natural” decomposition for many complex systems: e.g., body-brain co-evolution [39, 64] and multi-agent learning [65] in the evolutionary robotics field, and expectation-maximization [66, 67] from the AI field. Their significant feature is that all the subcomponents interact nonlinearly but work at different time scales or computing units (leading to different update frequencies). Note that the so-called ‘natural’ decomposition reflects the design preference or mandatory requirement (not necessarily the ‘optimal’ decomposition solution).
  • 2.
    In parallel/distributed computing, a single computing unit (e.g., a CPU core) always has a relatively limited memory and capability with Moore Law’s ending [68]. For big-data driven LSO with certain structures [69, 70, 71], general decomposition strategies (e.g., CD) are a viable solution for scalability, one critical metric for any general-purpose optimizer.
  • 3.
    There may be a relatively weak interaction between subcomponents in some (not all) non-separable real-world optimization problems, where decomposition-based methods could converge fast (within a practically accepted accuracy). This may partly explain why there always have some researchers and practitioners to use them since the establishment of the optimization area (furthermore, another advantage is their relative ease to understand and implement) [59, 71].
Convergence Analyses: In order to make it mathematically tractable, we simplify CC as a continuous-game (CG) model [72] but without losing its essential property. As compared to previous evolutionary game theory (EGT) [65, 73], our CG model can provide a much clearer analytical perspective built on only the pure Nash equilibrium (PNE) concept, much simpler than its mixed counterparts used in EGT. We theoretically show that under what conditions CC convergences, which most convex-quadratic benchmark functions can satisfy (see Section 3). To validate its prediction ability, we further demonstrate that many overlapping functions from the two newest test suites could essentially satisfy these conditions, which illustrates that CC finally converges to the global optimum on them, which is proved for the first time (our simulation experiments can perfectly match our predictions). Furthermore, using this new model, we reconfirm previously discovered pathologies of CC in a unified manner (i.e., relative generalization [74] and loss of gradients [75]). Depending upon these theoretical analyses, we propose to use a hierarchical decomposition structure to alleviate these issues and to obtain better generalizability (see Section 4.A).
Distributed Acceleration: Intuitively, CC appears to fit for distributed computing well [19]. However, under nonlinear inter-dependencies between subcomponents, the parallelism of CC is still a non-trivial task, since such a nonlinearity can lead to a difficulty in credit assignment for fitness evaluations of each subpopulation. To bypass this, we use the recently proposed multi-level learning/evolution (MLE) framework [76, 77], where each subpopulation conducts (local) metric learning on only its corresponding subspace for a much lower time and space complexity, which is important to match the hierarchical memory structure of CPU. At the same time, multiple large-scale CMA-ES variants (e.g., LM-CMA) are maintained in parallel, in order to reduce the possible risk of getting trapped into a suboptimal Nash equilibrium often encountered by CC. After each relatively short learning period, all learnt information will be collected, selected, and diversified at the meta-level for next cycles. Overall, the MLE framework can combine CC’s fine-tuning ability from decomposition with CMA-ES’s invariance [78] property (see Section 4).
Numerical experiments on a set of high-dimensional benchmark functions validate its search performance and scalability on an industry-level clustering computing platform with 400 cores (see Section 5).
2 Related Works
In this section, we first analyze the state-of-the-art CC and then discuss its real-world applications with non-separable forms and related theoretical advances. Next, we check many of partially separable real-world problems from the mathematical optimization community, to show that nearly all of them are non-separable according to variable interactions. Finally, following the suggestion from the latest CC survey [24], we build a connection between CC and its gradient-based counterpart (i.e., coordinate descent, CD) to help better understand the essence of decomposition-based optimizers.
2.1 State-of-the-Art CC
Since the publish of [32], a series of different improvements based mainly on DG (e.g., [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98]) have been proposed till now. Although these improvements often reduce the needed number of function evaluations, they might become over-skilled because they are built on only the PAS assumption [^6] The PAS assumption may be of a theoretical interest for some researchers.. As stated before, what kinds of real-world applications satisfy the PAS assumption is still an open question. Similar issue has happened in GA’s building- block hypothesis (BBH) for crossover operators (refer to [29, 99]). Like BBH, while PAS is intuitively appealing, finding functions from real-world applications to support it appeared surprisingly difficult. Although these automatic separability detection techniques sometimes could be used to analyze the variables interaction matrix as the basis of possible problem transformation for gray-box optimization [22], such a problem transformation is unavailable in box-box scenarios. Note that the real-world problem itself in [22] is non-separable (after transformation its new form is still non-separable). Similarly, in both [89] and [98], their tested real-world problems are also non-separable, despite they focused on decomposition.
Till now, the decomposition of non-separable problems is still a challenge. The above separability detection will become of less importance for non-separable functions when such a prior knowledge (i.e., non-separability) is relatively easy to obtain in practice (e.g., [100, 51] to name but a few). Recently, Chen et al. [101] considered non-additively partially separable problems [102]. However, how to extend it for general non-separable problems is still unclear. Komarnicki et al. [103] suffered from the same issue. Zhang et al. [104] applied CC to non-separable problems and obtained promising results on some benchmark functions. However, they did not analyze its convergence property. Jia et al. [105] extended contribution-based CC to a special type of non- separable functions (called overlapping [62, 106, 36]). Similarly, they did not analyze whether CC converges or not yet. Although recently Ge et al. [86] presented a simple convergence analysis under the block separability assumption, it is not suitable for non-separability.
2.2 Theoretical Advances of CC
There have been a series of theoretical works based on EGT to analyze complex convergence behaviors of CC on non-separable functions. Wiegand [74] for the first time proved that its replicator equations converge to PNE and showed that CC sometimes suffers from the relative generalization issue. Although the following-up works (e.g., [73, 107, 108]) extended Wiegand’s work somewhat, they are not extended to current large-scale CC versions due to the following factors: 1) It considers only one very simple decomposition case on a (discretized) low-dimensional search space; 2) it depends on a highly simplified payoff matrix for computing mixed Nash equilibria and therefore it cannot consider complex fitness landscape explicitly; 3) its derived lenient learning is still rarely used for LSO [65]. Overall, although they could provide a valuable analytical tool for CC, new theoretical advances are still expected for LSO-focused CC.
In [109] and [110], Jansen and Wiegand provided the first expected runtime analysis for CC. Surprisingly, they found that “the property of separability is neither a sufficient one to imply an advantage of CC, nor is inseparability a sufficient enough property to imply a disadvantage”. Similarly, Gomez et al. [51] argued that “much of the motivation for using the CC approach is based on the intuition that many problems may be decomposable into weakly coupled low-dimensional subspaces that can be searched semi-independently by separate species. Our experience shows that there may be another, complementary, explanation as to why cooperative coevolution in many cases outperforms single-population algorithms”. Overall, a deeper convergence analysis is highly desirable for recent large-scale CC versions.
2.3 Partially Separability and Coordinate Descent
Originally, the concept of partially separability (PS) [111] was defined in 1981 from the mathematical optimization field and is still studied by its one original author Toint now [112]. Until 2010, its very special form (i.e., PAS) was popularized for CC in the large-scale BBO field via a series of IEEE-LSGO competitions, despite there was one earlier (1999) EA paper also involving it [33]. Here, we re-check PS based on the original paper [^7] The first (1981) paper of partially separability is not accessible online. Unfortunately, the original author (Toint) cannot provide its pdf version (via email communication). Instead, we use his 1983 version, which provides a set of 50 functions. Again, this 1983 paper also becomes inaccessible online now (to our knowledge). Luckily, we have saved it locally before (if you want to read it, please contact Toint or us only for academic purpose).. Surprisingly, there is only one function in [111] that meets the PAS assumption in all 43 functions (here we have excluded 7 extra low-dimensional functions, because they are only used to test programming correctness). More et al. [113] collected a total of 24 optimization problems from many real-world applications. All [^8] We have excluded 4 problems since their objective function formula are not given explicitly, though they seems to be also non-separable. of them have a non-separable form [114]. Arguably, most PS instances from the mathematical optimization field are non-separable, obviously different from previous IEEE-LSGO competitions. We notice that some of recent CC papers start to extend previous IEEE-LSGO competitions to non-separable cases via e.g., overlapping functions. Similarly, coordinate descent (CD) is commonly used to optimize non-separable problems (often with friendly structures), since its first application in 1954 [115]. Note that for CD separability is limited to only the regularization term (in other words, the whole function form is generally inseparable).
3 Convergence Analyses via a Continuous-Game Model
In this section, we propose a continuous-game model as a basis of convergence analyses of CC. To validate its prediction ability, the convergence behaviors of CC are analyzed on common fitness models and the latest overlapping functions.
3.1 Continuous-Game (CG) Model of CC
For any objective function [^9] Only the minimization of objective (fitness) functions is considered, since maximization can be transformed into minimization simply by negating it. $f(\mathbf{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, CC divides all its coordinate indexes $\{1,\ldots,n\}$ into a set of mutually exclusive partitioning groups $p=\left\{g_{1},\ldots,g_{m}\right\}$ where $1<m\leq n$ and $g_{i}\neq\emptyset,g_{i}\cap g_{j\neq i}=\emptyset,\cup_{i}g_{i}=\{1,\ldots,n\},\forall i,j\in\{1,\ldots,m\}$. Note that $1<m$ excludes no decomposition case. An interesting theoretical question is how many possible partitions exist for decomposition-based methods. The Bell number, a computing concept from combinatorial mathematics, can help answer this question. For example, even for a 25-d function, there are totally 4,638,590,332,229,999,352 partitioning solutions (even when the partitioning order is not considered here) [116].
3.1.1 Ubiquity of Pure Nash Equilibrium (PNE)
In game theory, it is well-known that there always exists at least one mixed (not necessarily pure) Nash equilibrium for any noncooperative game [117, 118]. However, our previous conference paper has mathematically shown that CG’s special payoff structure guarantees the ubiquity of the pure Nash equilibrium (PNE) for any one partitioning (refer to [119] for details). Hereinafter, in order to avoid ambiguity, we will focus on only the PNE as defined below.
Definition (Pure Nash Equilibrium, PNE). Given any partition $p=\left\{g_{1},\ldots,g_{m}\right\}$ of the fitness function $f(\mathbf{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, we say that its decision vector $\mathbf{x}\in\mathbb{R}^{n}$ is one pure Nash equilibrium (w.r.t. $p$ ) iif the decision subvector $\mathbf{x}_{g_{i}}\in\mathbb{R}^{\left|g_{i}\right|}$ for each $g_{i}\in$ $p,\forall i\in\{1,\ldots,m\}$, satisfies
$f\left(\mathbf{x}_{g_{i}},\mathbf{x}_{\neq g_{i}}\right)\leq f\left(\mathbf{x}_{g_{i}}^{\sim},\mathbf{x}_{\neq g_{i}}\right),\forall\mathbf{x}_{g_{i}}^{\sim}\in\mathbb{R}^{\left|g_{i}\right|}\backslash\left\{\mathbf{x}_{g_{i}}\right\},$ (1)
where $\mathbf{x}_{\neq g_{i}}$ denotes all the remaining decision subvectors excluding $\mathbf{x}_{g_{i}}$ and $\mathbf{x}_{g_{i}}^{\sim}$ denotes any other possible values (i.e., $\left.\mathbb{R}^{\left|g_{i}\right|}\backslash\left\{\mathbf{x}_{g_{i}}\right\}\right)$. If (1) is strict, we say that $\mathbf{x}$ is a strict PNE (w.r.t. $p$ ). Since this concept relies heavily on the pre-partitioning, we add a suffix “(w.r.t. $p$)” to emphasize and clarify it, if necessary. To help understand this solution concept for decomposition-based optimization, we visualize its distributions on four common non-separable test functions in Fig. 1.
(a) $f_{1}(x,y)=7x^{2}+6xy+8y^{2}$
(b) $f_{2}(x,y)=x^{2}+10y^{2}$ (rotated)
(c) $f_{3}(x,y)=100(x^{2}-y)+(x-1)^{2}$
(d) $f_{4}(x,y)=|x-y|-\min(x,y)$
Figure 1: Visualization of all pure Nash equilibria (PNE) on four 2-d non-separable functions. The white line represents the contour levels of fitness landscape while the green point/line (aka best-response curve) denotes the location(s) of PNE.
As is seen in Fig. 1, the top-left is a strictly-convex function, designed by Shi et al. [71] from the latest CD survey. There is a unique PNE, which is the global optimum (CC could converge to it rapidly owing to weak dependency). The top-right is an ill-conditioned (rotated) Ellipsoid function commonly used to benchmark the local search ability of EAs. Similarly, there is only one PNE (CC still could converge to it but its convergence rate is very slow owing to strong dependency). The bottom-left is a very famous non-convex function [120], where there is also a unique PNE (CC still could converge to it but the convergence rate is slow when approaching the parabolic curve). The bottom-right is a convex but non-differentiable function, originally proposed by Warga [121] from the CD community. For it, there is a continuum of PNE (CC does not necessarily converge to the global optimum but to other suboptimal PNE, depending on the starting point). For their convergence demonstrations, please refer to B.
Lemma 1: Given any partitioning solution $p=\left\{g_{1},\ldots,g_{m}\right\}$ of $f(\mathbf{x})$, its global optimum $\mathbf{x}^{*}$ is a pure Nash equilibrium w.r.t. $p$. However, the inverse is not necessarily true.
The ubiquity of PNE under any partition can be guaranteed by the above lemma, only if the objective function has (at least) one global optimum (which is a default assumption in most theoretical analyses).
Lemma 2: Given any partition $p=\left\{g_{1},\ldots,g_{m}\right\}$ of $f(\mathbf{x})$ : $\mathbb{R}^{n>2}\rightarrow\mathbb{R}$ where $\left|g_{i}\right|>1,\exists i\in\{1,\ldots,m\}$, we can divide $p$ along $g_{i}$ and get a new partition $p^{\prime}$ satisfying $\left|p^{\prime}\right|>|p|$. If $x$ is a PNE w.r.t. $p$, then $\mathbf{x}$ is also a PNE w.r.t. $p^{\prime}$. However, the inverse does not necessarily hold. (We term this interesting property as the downward rather upward propagation of $PNE$.)
Since lemma 2 plays a fundamental role in the following convergence analyses, we design a simple example to illustrate it. Consider an artificially designed function case $f(x,y,z)=$ $f(x,y)+f(y,z)$, where $f(x,y)=(x+y)^{2}+(y-1)^{2}$ and $f(y,z)=(y+1)^{2}+(y+z)^{2}$. For it, there are 4 partitions: $\{\{1,2\},\{3\}\},\{\{1,3\},\{2\}\},\{\{2,3\},\{1\}\}$, and $\{\{1\},\{2\},\{3\}\}$, as shown in Fig. 2 (for simplicity, the order of partitions is not considered here). Interestingly, all partitions can be well reorganized in a hierarchical (recursive) manner. Each green arrow represents one further partition operation (i.e., from $p$ to $p^{\prime}$ ). The global optima, a special type of PNE, can propagate both downwardly (from $p$ to $p^{\prime}$ ) and upwardly (from $p^{\prime}$ to $p$ ). One theoretical significance of lemma 2 is that it can greatly simplify the difficult seeking process of PNE, as proved in the next subsection.
Figure 2: A simple example to illustrate the downward rather upward propagation of PNE along hierarchical partitions.
3.1.2 Convergence to a PNE in CG
Although it is well-known in the scientific community that “essentially all models are wrong[73], a good model for CC should meet (at least) the following two criteria while omitting the peripheral details: 1) It should still capture the essence of the algorithm even after simplifications, which can be used to predict the limit convergence behavior; 2) It should exhibit somewhat prediction and even extrapolation ability, despite it is nearly impossible to exclude the failure risk on some unseen practical scenarios.
In previous EGT models [73, 74], CC was modeled using two-player replicator equations. Alternatively, CC can also be simply modeled as a continuous game (CG) on the original continuous search space, which is comprised of multiple dynamically-coupled subpopulations. Continuous games have been investigated in many research areas (economics [122], nonlinear automatic control [123, 72], multi-agent learning [124], engineering design [125], and AI [126, 127], etc.).
In principle, our CG model can be seen as a particular form of CG proposed by Ratliff et al. [72] from the automatic control community. For the former, all the players have the same objective function form but with a disjoint subset of decision vectors. However, for the latter, each player has its own objective function, which generally has a different form with each other. We borrow the classical model [128, 129] from CD to formalize each cycle of CC, mathematically represented as one transformation $T:\Omega\subset\mathbb{R}^{n}\rightarrow\Omega\subset\mathbb{R}^{n}$ of the search space $\Omega$ into $\Omega$ itself, satisfying the two following conditions:
  1. 1.
    $f(T(\mathbf{x}))\leq f(\mathbf{x})$, where $\mathbf{x}$ is the best-so-far solution;
  2. 2.
    $f(T(\mathbf{x}))=f(\mathbf{x})\Leftrightarrow T(\mathbf{x})=\mathbf{x}$. (a fixed point)
Condition 2 excludes one very special case (that is, when $f(T(x))=f(\mathbf{x}),\mathbf{T}(\mathbf{x})\neq\mathbf{x})$, which otherwise results in the cyclical behavior of the best-so-far solution and prevents the global convergence (which was found first by Powell in 1973 [130]).
For successive transformation processes, the key point is to extract a strictly decreasing fitness subsequence from the nonincreasing fitness sequence of the best-so-far solution $\mathbf{x}$ generated by $\mathrm{CC}$. This could be guaranteed by a strong assumption that each suboptimizer has sufficient search ability for each subproblem (that is, it can find the global minimizer for each subproblem given a finite number of iterations). It is worthwhile noting that even such a strong assumption cannot ensure the convergence to the global optimum on the original problem (see Fig. 1 for an example). As a result, finally $\lim_{t\rightarrow+\infty}T_{t}(\mathbf{x})$ will move towards a fixed point (that is, ${T}(\mathbf{x})=\mathbf{x})$, which is also a PNE [^10] Its mathematical proof is simple via proof by contradiction: In fact, the Nobel-prize winner Nash was the first to connect the fixed point with the equilibrium point (now named after him) in his seminar PNAS and Annals of Mathematics papers, laying the theoretical foundation of modern game theory.. Please refer to A for a detailed mathematical demonstration. In practice, the global minimizer of each subproblem in each cycle is not needed to find exactly [62] in each cycle.
We acknowledge that the above CG model is very simple from a practical view of point. Based on it, however, we will derive some interesting theoretical results, as presented below.
3.2 Convergence Analyses on Convex Functions
Based on the above CG model, the convergence problem of CC can be well converted to the seeking and classification problem of PNE. Here we provide a main theorem on convex functions, as presented below:
Main Theorem: Consider any partitioning solution $p=\left\{g_{1},\ldots,g_{m}\right\}$ of a continuous-differentiable convex function $f(\mathbf{x})\in{C}^{1}(\Omega,\mathbb{R})$, where $\Omega$ is an open convex set [^11] Here it is implicitly assumed that there is (at least) one global optimum in this open convex set. in $\mathbb{R}^{n}$. For $f(\mathbf{x})$, all PNE (w.r.t. $p$) are equivalent to the global optima, and vice versa.
Proof: First we demonstrate one special partition case $p^{n}=$$\{\{1\},\ldots,\{n\}\}$ (i.e., $m=n$ ). Assume that $\widetilde{\mathbf{x}}$ is a PNE (w.r.t. $p^{n}$). By the definition of PNE, we can simply infer that $\widetilde{\mathbf{x}}_{\{i\}}\in$ $\underset{x_{\{i\}}}{{argmin}}f\left(\mathbf{x}_{\{i\}},\mathbf{x}_{\neq\{i\}}\right)\subseteq\Omega,\forall i\in\{1,\ldots,n\}$.
Owing to the continuous differentiability of $f(x)$, each partial derivative $\nabla_{\{i\}}f(\widetilde{x})=0,\forall i\in\{1,\ldots,n\}$. So, $\tilde{x}$ is also a stationary point. Note that for $\underset{x_{(i)}}{{argmin}}f\left(\mathbf{x}_{\{i\}},\mathbf{x}_{\neq\{i\}}\right)$, the solution is not necessarily unique. According to the convexity of $f(\mathbf{x})$, it is well-known that all stationary points are the global optima. With lemma 1 (that is, all global optima are PNE), we conclude that for this special case $p^{n}=\{\{1\},\ldots,\{n\}\}$, all PNE (w.r.t. $p^{n}$) are the global optima, and vice versa.
Then, we show the general case for any other partition $p=$$\left\{g_{1},\ldots,g_{m}\right\}$ when $m\neq n$. Based on downward propagation of PNE (lemma 2), all PNE w.r.t. $p$ belongs to a set of PNE w.r.t. $p^{n}$, which are also global optima. With lemma 1 , we finish the proof perfectly.
Clearly the above main theorem is valuable to understand why CC could converge to the global optima on many (though not all) non-separable functions. In effect, many of common test functions fall into this interesting class [131]. Take e.g., Cigar, Discus, CigarDiscus, Ellipsoid, and Schwefel as examples for convex-quadratic functions [5]. Note that their non-separable rotated-and-shifted versions still fall into this class. Surprisingly, all the overlapping functions from the newest LSO test suite for CC [62] essentially still meet these above assumptions (after suitable simplifications). The above main theorem can in part explain why one state-of-the-art CC version could obtain very competitive performance on a particular class of overlapping functions, despite all of them are non-separable (refer to the following subsection for more details). In most of previous CC papers for LSO, both the theoretical and practical significance of PNE are overlooked or blurred. However, in the non-separable cases, there are more complex relationships between PNE and global optima, which need to be clarified theoretically [132].
3.3 Global Convergence Guarantee on Overlapping Functions
In order to validate the extrapolation ability of our proposed CG model, here we consider a class of overlapping (a special type of non-separability) functions from the latest LSO test suite [62]. The original experiments reported in [62] showed that on each function the best-so-far solution is far from the global optimum, since a relatively small maximum of function evaluations (i.e., 3e6) was used. Based upon corollary 1 shown below, however, we predict that CC could converge to the global optimum on them finally, since they essentially meet the assumptions of the main theorem, when two benchmarking operations (i.e., non-smoothing and asymmetry) are removed to make the involved mathematics tractable. To validate our prediction, we increase the maximum of function evaluations from 3e6 to 3e7 significantly. New experiment results are in accordance with our theoretical prediction [^12] The source code is freely available at https://bitbucket.org/yuans/rdg3.(even when two benchmarking operations are added).
Corollary 1: Consider a non-separable overlapping function $f(\mathbf{x})=$ $\sum_{i=1}^{m}f_{i}\left(\mathbf{x}_{i}\right)$ with $\mathbf{x}\in\Omega\subseteq\mathbb{R}^{n},\mathbf{x}_{i}\cap\mathbf{x}_{j}\neq\emptyset,\exists i\neq j\in$$\{1,\ldots,m\}$ and $U_{i}\mathbf{x}_{i}=\mathbf{x}$. All subfunctions $f_{i}$ have the exactly same [^13] In effect, such a condition can be further weakened, only if it is convex and continuous-differentiable. For simplicity, however, we exclude it here. base form $f_{b}(\mathbf{x})=\sum_{i=1}^{n}\left(\sum_{j=1}^{i}\mathbf{x}_{i}\right)^{2}$ (i.e., Schwefel’s Problem 1.2 in [62]). Its rotated-and-shifted version $f_{s}^{r}(\mathbf{x})=$ $f_{b}(\mathbf{R}(\mathbf{x}-\mathbf{s}))$ is used by each $f_{i}\left(\mathbf{x}_{\mathbf{i}}\right)$, where $\mathbf{R}$ is an orthogonal matrix and $s$ is a shift vector. For this class of overlapping functions, all PNE (w.r.t. any partition) are global optima, and vice versa.
Proof: First we show the each continuous-differentiable subfunction $f_{i},\forall i\in\{1,\ldots,m\}$, is convex. Since all $f_{i}$ share the same form $f_{b}(\mathbf{x})=\sum_{i=1}^{n}\left(\sum_{j=1}^{i}\mathbf{x}_{j}\right)^{2}$, we only need to prove the base form case. Its Hessian $\mathbf{H}$ is
$2*\left[\begin{array}[]{cccccc}n&n-1&n-2&&2&1\\ n-1&n-1&n-2&\cdots&2&1\\ n-2&n-2&n-2&&2&1\\ &\vdots&&\ddots&\vdots\\ 2&2&2&\cdots&2&1\\ 1&1&1&&1&1\end{array}\right]$
Successively adding a multiple of row $i-1$ to another row $i$ by $-\frac{n-(i-1)}{n-(i-2)},i=n,n-1,n-2,\ldots,3,2$, we get a new matrix:
$\mathbf{A}=2*\left[\begin{array}[]{cccccc}n&n-1&n-2&&2&1\\ 0&a_{22}&a_{23}&\cdots&a_{2(n-1)}&a_{2n}\\ 0&0&a_{33}&&a_{3(n-1)}&a_{3n}\\ &\vdots&&\ddots&\vdots&\\ 0&0&0&\ldots&a_{(n-1)(n-1)}&a_{(n-1)n}\\ 0&0&0&&0&1/2\end{array}\right],$
where $a_{ii}=(n-(i-1))-\frac{n-(i-1)}{n-(i-2)}*(n-(i-1))>0,i=$ $n,n-1,n-2,\ldots,3,2$. By the theorems from linear algebra (pp. 60-61 in [133]), $\mathbf{H}$ is row equivalent to $\mathbf{A}$. With the theorem from linear algebra (pp. 370 in [133]), we conclude that $\mathbf{H}$ is a positive definite matrix, which means that $f_{b}(\mathbf{x})$ is a (strictly) convex function.
Second, we consider its rotated-and-shifted variant $f_{s}^{r}(\mathbf{x})=$ $f_{b}(\mathbf{R}(\mathbf{x}-\mathbf{s}))$, where $\mathbf{R}(\mathbf{x}-\mathbf{s})$ is still a convex set. For $\forall\mathbf{x}\neq$ $\mathbf{y}\in\Omega$ and $\theta\in[0,1],f_{s}^{r}(\theta\mathbf{x}+(1-\theta)\mathbf{y})=f_{b}(\mathbf{R}(\theta\mathbf{x}+$ $(1-\theta)\mathbf{y})-\mathbf{s}))=f_{b}(\theta\mathbf{R}(\mathbf{x}-\mathbf{s})+(1-\theta)\mathbf{R}(\mathbf{y}-\mathbf{s}))$, by the convexity of $f_{b},\leq\theta f_{b}(\mathbf{R}(\mathbf{x}-\mathbf{s}))+(1-\theta)f_{b}(\mathbf{R}(\mathbf{y}-\mathbf{s}))=\theta f_{s}^{r}(\mathbf{x})+(1-\theta)f_{s}^{r}(\mathbf{y})$. Therefore, its rotated-and-shifted variant is still a convex function.
Then, if $f_{i}(\mathbf{x}),\forall i=1\ldots,m$, is a convex function, their nonnegative weighted sum is still a convex function (pp. 79 in [134]). A total of 20 overlapping functions in [62], no matter with conforming or conflicting components, still meet it. With the main theorem, we can conclude the proof.
Considering that recent $\mathrm{CC}$ works focus on the non-separable overlapping cases, the above corollary can provide a relatively deep theoretical understanding to CC’s convergence properties on some overlapping functions with certain structures. Interestingly, the above corollary can also be extended to some overlapping functions designed in [105], since their base form is also convex.
4 Distributed Multi-level Learning
In this section, a distributed cooperative coevolution (DCC) framework is proposed where the recent multilevel learning [77, 76] is employed, in order to improve the efficiency and effectiveness of CC on modern clustering computing platforms. Given the high complexity of distributed algorithms, we provide its source code available at GitHub [^14] https://github.com/Evolutionary-Intelligence/DCC to ensure repeatability. For distributed algorithms, multiple performance trade-offs must be considered: such as, distributed scheduling, distributed (shared) memory management, network communications, fault tolerant, system control, and so on. In this paper, we focus on the application-level parallelism and leave other tedious tasks to the underlying distributed computing engine.
4.1 Hierarchical Structure of DCC
Algorithm 1 Distributed Cooperative Co-evolution (DCC) with CMA-based Multilevel Learning. (Refer to its open-source Python code for repeatability: https://github.com/Evolutionary-Intelligence/DCC.)
$p$: total number of available slave nodes in computing platform
1:$p\_es$: number of slave nodes to run LM-CMA
[…]
3:$b\_f$: the best-so-far fitness
4:Initialize: $b\_x\leftarrow\infty,b\_y\leftarrow\infty$
[…]
6:use Meta-ES to set global step-size
7:use elitist or weighted averaging to set distribution mean
[…]
12:update distribution mean via weighted averaging at meta-level
13:conduct collective learning of covariance matrices $\triangleright$ meta-diversity
[…]
As previously shown, CC tends to convergence to the PNE. If the PNE is not the global optimum (refer to Fig. 1 as an example), CC easily gets trapped into this suboptimal solution. This issue (known as relative generalization) is attributed to decomposition, which cannot be avoided in essence but may be alleviated somewhat. On a class of ill-conditioned non-separable functions [79], like its gradient-based counterpart (CD), CC often shows much slower convergence rates than second-order-type optimizers (e.g., LM-CMA and L-BFGS). However, when as a general-purpose BBO framework for LSO, it is highly desirable that CC could also handle these challenging issues. To achieve this goal, we use the recently proposed multilevel learning framework [77, 76] to combine the best of both worlds (i.e., the powerful invariance property of LM-CMA and the fine-tuning ability of CC via CMA-ES on low-dimensional subspace). In this paper, we only consider the hierarchical structure to implement it for simplicity, as seen in Fig. 3. In principle, a more general recursive structure can be also applied here, which we leave for a future work.
Figure 3: A hierarchical structure of our distributed cooperative coevolution framework (DCC) for large-scale BBO.
As shown in Fig. 3, there are multiple slave nodes as well as one master node for distributed computing (e.g., on Spark [135] and Ray [136]). For DCC, at the meta-level the master node mainly coordinates different optimizers, each of which is run in one separate computing unit (a CPU core). Here we choose two different ES versions (CMA-ES [27] and LM-CMA [60]) as two search engines for the low-dimensional (often $<50$) subspace (after decomposition) and the original large-scale ($\gg 100$) space, respectively. In principle, any other EAs (e.g., GA [137], EP [138], PSO [139], DE [140], and EDA [141]) can be also selected as the suboptimizer. The rationale behind using two search engines lies that in the original space LM-CMA has a high possibility to escape from the suboptimal PNE via learning of promising evolution paths while maintaining a low computationally complexity (see [60] for analysis) and in the decomposed subspace the full-fledged CMA-ES [27] can well approximate the inverse of its Hessian matrix for fine-tuning with a reasonable memory consumption on each single computing unit (which is vital for scalability of distributed algorithms for LSO). Refer to the following three subsections as well as Algorithm 1 for more details.
4.2 Collective Learning of Covariance Matrices
For both the LM-CMA and the CMA-ES, covariance matrix adaptation (CMA) plays a central role in their invariance property. Owing to its cubic (or quadratic after modifications) time complexity, CMA is difficult to directly apply to LSO. Instead, approximating CMA with low-memory or low-rank techniques is more acceptable for LSO, resulting in lower time and space complexities (e.g., $O(n*\log(n))$ in [60]). For instance, CMA’s rank-one update (without the rank-$\mu$ update) can be finally transformed to the following nonrecursive form for offspring sampling (refer to [142] for a detailed mathematical deduction):
$\mathbf{d}^{t}=$ $\left(\left(1-\frac{c_{1}}{2}\right)\mathbf{I}+\frac{c_{1}}{2}\mathbf{p}^{1}\left(\mathbf{p}^{1}\right)^{T}\right)*\ldots$
$*\left(\left(1-\frac{c_{1}}{2}\right)\mathbf{I}+\frac{c_{1}}{2}\mathbf{p}^{t-1}\left(\mathbf{p}^{t-1}\right)^{T}\right)$
$*\left(\left(1-\frac{c_{1}}{2}\right)\mathbf{I}+\frac{c_{1}}{2}\mathbf{p}^{t}\left(\mathbf{p}^{t}\right)^{T}\right)\mathbf{z}^{t},$
where $\mathbf{d}^{t}\in\mathbb{R}^{n},\mathbf{I}\in\mathbb{R}^{n*n},\mathbf{p}^{t}\in\mathbb{R}^{n},\mathbf{z}^{t}\in\mathbb{R}^{n}\sim N(\mathbf{0},\mathbf{I}),0<c_{1}$ $<1$ are the directional vector, identity matrix, evolution path, standard Gaussian permutation, and learning rate of the rank-one update for iteration $t$, respectively. In the actual code implementation, $\mathbf{I}$ is omitted since $\mathbf{I}\mathbf{z}^{t}=\mathbf{z}^{t}$ The above form should be implemented from right to left, in order to avoid the too expensive matrix operation. Only $m=O(\log(n))\ll n$ dot products (with $O(n*\log(n))$ computational complexity) are involved when sampling each offspring.
Although they fit for distributed computing owing to their low memory requirements, these approximations may lose sufficient diversity on the covariance matrix: 1) the rank-$\mu$ update is not used, which is beneficial for rugged fitness landscapes; 2) the rank-one update is mainly beneficial for a predominated search direction (e.g., Cigar and Rosenbrock [5]) rather than multiple promising search directions (e.g., Discus and Ellipsoid [5]); 3) the exponentially smoothing update exploits mainly the temporal information but not the (parallel) spatial information, which can be alleviated via the distributed algorithms like Meta-ES [29].
In this subsection, we use the collective learning (or called collective intelligence by Schwefel [143]) paradigm to exploit both the spatial and temporal information for CMA, since this paradigm can naturally match distributed computing. On each slave node, its corresponding ES learns different covariance matrices (directional vectors) periodically. Different form LM-CMA, for all the CMA-ES only the low-dimensional subspace is searched, which needs an extra assembling operation in the optimizer-level of CC (see Fig. 3). After each relatively short learning period, all information is collected into the meta-level of DCC. Since the temporal information has been exploited on the slave nodes, we average all the covariance matrices of the better LM-CMA instances (e.g., fixed to 1/5 of all instances) as its base of the next cycle. The (weighted) averaging does not involve the expensive matrix multiplication operation at the meta-level, which otherwise can severely degrade the speedup on mainstream distributed computing systems based on the master-slave architecture (according to the well-known Amdahl’s Law). For better diversity of covariance matrix at the meta-level, we assemble a set of direction vectors from LM-CMA with covariance matrices from CMA-ES for some new LM-CMA instances in the next cycle.
Overall, it is expected to maintain the powerful invariance property of LM-CMA especially for ill-conditioned landscapes (really a challenge for CC), while keeping the fine-tuning ability of CC over a cycle of different subspaces. Our theoretical results have shown that CC tends to converge to the global optimum under certain conditions (even given any decomposition).
4.3 Distributed Meta-ES for Global Step-Size Adaptation
In the serial CMA-ES [28], the global step-size adaptation is decoupled with the adaptation of covariance matrix, since they can work independently at different time scales for better convergence progress. In the standard CMA-ES form, the cumulative step-size adaptation (CSA) based on the evolution path is used to set the global step-size on-the-fly, which exploits the temporal correlation information over successive generations to speed up convergence significantly. However, in the distributed computing context, the CSA can be further improved, since the spatial information becomes accessible in parallel, resulting in the Meta-ES [29]. Following our previous work [76], we combine Meta-ES with CSA to enjoy the best of both worlds. Specifically, Meta-ES is run at the meta-level of DCC to keep spatial diversity while CSA is employed for each serial ES instance to exploit temporal correlation.
4.4 Distribution Mean Update via Weighted Averaging
The mean update of the normal search distribution has a significant impact on both the stability and effectiveness of DCC. Following the theoretical work from Beyer [29], we use the (weighted) averaging strategy at the meta-level, in order to obtain the genetic repair effect, which is also consistent with all ESs in slave nodes. Note that this simple strategy is also used in some gradient-based distributed optimizers (e.g., for DNN training). However, we have previously observed in [76] that sometimes it could result in the regression issue on some functions, which was also identified by Rudolph in [144]. To stabilize the evolution process and keep diversity at the meta-level, we also keep maintaining a proportion of elitist LM-CMA instances in parallel at each learning period (e.g., fixed to 1/20, typically depending on the number of available CPU cores).
In summary, we use a state-of-the-art clustering computing system called Ray [136], developed mainly by a research team from UC Berkeley, to implement our distributed algorithm (DCC).
5 Numerical Experiments
In this section, large-scale numerical experiments on a set of high-dimensional test functions are conducted, in order to show the advantages (and possible disadvantages) of DCC for large-scale black-box optimization.
5.1 Benchmarking Algorithms
We choose a total of 43 benchmarking baselines, which are classified into the following 10 main families: CC, ES, natural evolution strategies, cross-entropy methods, estimation of distribution algorithms, differential evolution, particle swarm optimization, genetic algorithms, simulated annealing, and random search [^15] Here we have excluded the classical evolutionary programming, since its modern versions for continuous optimization is very similar to ES.. For each algorithmic family, there are some standard versions and the latest large-scale variants. All their source code is taken from a recently developed open-source library for population-based optimization (called PyPop7161616https://github.com/Evolutionary-Intelligence/pypop), which is actively maintained now. Owing to page limits, please refer to https://pypop.readthedocs.io for their references.
5.2 High-Dimensional Test Functions
Table 1: A Set of 10 Benchmarking Functions
Function Name Mathematic Function
sphere $f(x)=\sum_{i=1}^{n}x_{i}^{2}$
cigar $f(x)=x_{1}^{2}+10^{6}\sum_{i=2}^{n}x_{i}^{2}$
discus $f(x)=10^{6}x_{1}^{2}+\sum_{i=2}^{n}x_{i}^{2}$
cigar__\__discus $f(x)=x_{1}^{2}+10^{4}\sum_{i=2}^{n-1}x_{i}^{2}+10^{6}x_{n}^{2}$
ellipsoid $f(x)=\sum_{i=1}^{n}10^{\frac{6(i-1)}{n-1}}x_{i}^{2}$
different__\__powers $f(x)=\sum_{i=1}^{n}\left|x_{i}\right|^{\frac{2+4(i-1)}{n-1}}$
schwefel221 $f(x)=\max(\left|x_{1}\right|,\cdots,\left|x_{n}\right|)$
step $f(x)=\sum_{i=1}^{n}(\lfloor x_{i}+0.5\rfloor)^{2}$
rosenbrock $f(x)=100\sum_{i=1}^{n-1}(x_{i}^{2}-x_{i+1})^{2}+\sum_{i=1}^{n-1}(x_{i}-1)^{2}$
schwefel12 $f(x)=\sum_{i=1}^{n}(\sum_{j=1}^{i}x_{j})^{2}$
Since the focus of this paper is non-separable large-scale BBO, we choose 10 high-dimensional test functions, which are often used to analyze the convergence property in the ES community. For modern ESs, invariance is one fundamental design principle when optimizing non-separable functions. See Table 1 for their detailed formula. As a standard benchmarking [145] practice, we use the rotation and shift operation to make the test function non-separable and avoid the origin being the global optimum, respectively.
In this paper, we set the number of dimensions to 2000 for all test functions. Since the rotation operation [^17] Generating rotation matrix via Schmidt orthogonalization has a cubic time complexity. has a quadratic computational complexity, we need to efficiently utilize the shared memory of each node to avoid expensive data-transfer operations, following our previous paper [76]. Note that this setting is critical for the speedup of memory-costly function evaluations for distributed EAs. For all test functions, the initial search range is set to $(-10.0,10.0)^{n}$.
5.3 Experimental Setups
On all test functions, each algorithm was totally run 5 times. We set two termination conditions for all algorithms to make fair comparisons as much as possible: 1) when the best-so-far fitness is below $1\mathrm{e}{-10}$; 2) when the maximum runtime exceeds 3 hours. As a result, the total CPU runtime roughly equals 6600 (= 44 algorithms*10 functions*5 trails*3 hours) hours (i.e., 275 days) if they are run only on a single machine. To reduce the time-consuming benchmarking process, 10 HPC servers were used in parallel, all of which were also together used to build a (slightly heterogeneous) clustering computing platform. There is a total of 400 CPU logic cores and about 340GB memory in our private clustering computing platform. Under page limits, please see online Supplementary Materials (available at https://github.com/Evolutionary-Intelligence/DCC) for detailed hardware and software configurations.
5.4 Experimental Results and Analyses
Figure 4: Median convergence curves on a set of 2000-d rotated and shifted test functions. To avoid crowd and confusion (given 44 algorithms), only the first three of top-ranked optimizers have a legend and only half (22) of optimizers are drawn in each subfigure. Refer to online Supplementary Materials for detailed data.
Figure 5: Median convergence curves on a set of 2000-d rotated and shifted test functions. To avoid crowd and confusion (given 44 algorithms), only the first three of top-ranked optimizers have a legend and only half (22) of optimizers are drawn in each subfigure. Refer to online Supplementary Materials for detailed data.
In Fig. 4 and 5, the convergence curves on all test functions are drawn for 22 black-box optimizers. To avoid crowd in plotting, all other 22 optimizers are shown in online Supplementary Materials (available at https://github.com/Evolutionary-Intelligence/DCC), since all of them are much worse than or close (only on one test function) to our distributed algorithm on these high-dimensional functions.
On five test functions (i.e., schwefel12, discus, ellipsoid, different__\__powers, and step), clearly DCC showed the best convergence rates. For the first forth cases, there are multiple (rather than one predominated) promising search directions. DCC’s diversity maintaining of covariance matrix can exploit them in parallel at the meta-level. For the last case, properly setting the global step size on-the-fly is important for fast convergence rate since there exist many plateaus. For DCC, its Meta-ES-based adaptation strategy can alleviate this problem significantly, since it exploits the spatial information well (at the same time multiple parallel CSA strategies could exploit the temporal information inside each cycle).
On four other functions (i.e., cigar, sphere, cigar__\__discus and rosenbrock), DCC still obtained very competitive results, though not the best one. For all of them except sphere there is one predominated search direction, which can be efficiently learnt by the (exponential smoothing) evolution path. In this case, keeping the diversity on covariance matrix seems to be not the most important. The performance differences on these functions are very small (mainly from the extra overhead of distributed computing), which may be negligible. For DCC, the stabilization strategy for distribution mean update helps to maintain the advantages of the underlying suboptimizers.
However, distributed computing is not a panacea, which cannot escape from the No-Free-Lunch theorems [132]. On schwefel221, DCC suffered from a slow convergence. This is due to the loss of gradient on this min-max function. See C for a theoretical analysis regarding convergence.
Overall, the zig-zag-type convergence rate is reminiscent of a famous natural evolution theory named as “punctuated equilibrium[146]. Such a punctuated equilibrium-style convergence curve may be an essential property of many distributed EAs, which is worthwhile to be further investigated.
6 Conclusion
In this paper, we have for the first time examined the coherence of the theoretical concept (PAS) for CC with many real-world applications, which is believed to be of central importance [113], and showed that few real-world applications could well match such an ideal assumption [^18] See https://tinyurl.com/3d4nnneb for a variety of applications regarding evolutionary computation.. Instead, we have provided a pure Nash equilibrium (PNE) perspective as a new avenue to capture the essence of CC [147] to explain currently empirical successes on some (rather all) non-separable LSO problems. Furthermore, based on the above theoretical results, we have extended the serial version of CC to the powerful distributed computing scenario and proposed its distributed version (DCC), where collective learning [77, 143] plays an important role when combined with two state-of-the-art ES variants (LM-CMA and CMA-ES). Numerical experiments have shown the scalability (w.r.t. CPU cores) and potentials of DCC on a set of high-dimensional benchmark functions. We plan to extend DCC in the future as follows: 1) to build a model to analyze its convergence rate; 2) to further validate its efficiency and effectiveness on some real-world applications; 3) to scale it up to more than one thousands of CPU cores (“more is different[148]).
Appendix A Convergence to a Pure Nash Equilibrium (PNE) for Cooperative Coevolution (CC)
Theorem 1: Given any partition $p=\left\{g_{1},\ldots,g_{m}\right\}$ of the objective function $f(\mathbf{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, where $1<m\leq$ $n$, we say that the convergence point of CC under this partition is also one pure Nash equilibrium (PNE).
Proof: If $\mathbf{x}^{*}=\left(x_{1}^{*},x_{2}^{*},\cdots,x_{n}^{*}\right)=\left(\mathbf{x}_{g_{1}}^{*},\cdots,\mathbf{x}_{g_{m}}^{*}\right)$ is a convergence point of CC under the partition $p=$$\left\{g_{1},\ldots,g_{m}\right\}$, then $\mathbf{x}_{g_{i}}^{*}$ is one of the global optima of the function $f\left(\mathbf{x}_{g_{1}}^{*},\cdots,\mathbf{x}_{g_{i}},\cdots,\mathbf{x}_{g_{m}}^{*}\right),i=1,\cdots,m$.
If $\mathbf{x}^{*}=\left(\mathbf{x}_{g_{1}}^{*},\cdots,\mathbf{x}_{g_{m}}^{*}\right)$ is not one pure Nash equilibrium (w.r.t. $p$ ), according to the definition of Pure Nash Equilibrium (PNE), we have that $\exists i\in\{1,\cdots,m\},\exists\check{\mathbf{x}}_{g_{i}}\in\mathbb{R}^{|g_{i}|}\backslash\left\{\mathbf{x}_{g_{i}}^{*}\right\}$, such that $f\left(\mathbf{x}_{g_{1}}^{*},\cdots,\check{\mathbf{x}}_{g_{i}},\cdots,\mathbf{x}_{g_{m}}^{*}\right)<f\left(\mathbf{x}_{g_{1}}^{*},\cdots,\mathbf{x}_{g_{i}}^{*},\cdots,\mathbf{x}_{g_{m}}^{*}\right)$, namely, then $\mathbf{x}_{g_{i}}^{*}$ is not the global optimum of $f\left(\mathbf{x}_{g_{1}}^{*},\cdots,\mathbf{x}_{g_{i}},\cdots,\mathbf{x}_{g_{m}}^{*}\right)$. It is a contradiction.
Note that we assume that for CC, the suboptimizer could obtain the global optimum for each subproblem given a limited number of function evaluations (but which could be any large number). We admit that such an assumption appears to be difficult to satisfy in practice. However, it helps to understand the convergence behavior of CC and capture the game-theoretical essence of CC.
Appendix B Convergence Analyses on Four Representative Test Functions
B.1 Function 1
$f_{1}(x,y)=7x^{2}+6xy+8y^{2}$
Since its Hessian matrix $\left[\begin{array}[]{ll}7&6\\ 6&8\end{array}\right]$ is positive definite, $f_{1}$ is differentiable strictly convex, then it has a unique global optimum $(0,0)$. Given $\left(x_{0},y_{0}\right)$ is a PNE, by the definition of PNE, $\left(x_{0},y_{0}\right)$ is the unique global optimum of differentiable strictly convex $f_{1}\left(x,y_{0}\right)$ and $f_{1}\left(x_{0},y\right)$. Then $\frac{\partial f_{1}\left(x,y_{0}\right)}{\partial x}=14x+6y_{0}=0$, we have $x=-\frac{6y_{0}}{14}=x_{0},\frac{\partial f_{1}\left(x_{0},y\right)}{\partial y}=16y+6x_{0}=0$, we have $y=$ $-\frac{6x_{0}}{16}=y_{0}$. So, $\left(x_{0},y_{0}\right)=(0,0)$.
B.2 Function 2
$f_{2}(x,y)=x^{2}+10^{6}y^{2}$
Since its Hessian matrix $\left[\begin{array}[]{cc}1&0\\ 0&10^{6}\end{array}\right]$ is positive definite, $f_{1}$ and its rotation variant are all strictly convex and have the unique global optimum $(0,0)$. Like the above proof, they have only one PNE, namely global optimum $(0,0)$.
B.3 Function 3
$f_{3}(x,y)=100\left(x^{2}-y\right)^{2}+(x-1)^{2}$
Obviously, $f_{3}\geq 0$ and
$\left\{\begin{array}[]{c}\frac{\partial f_{3}(x,y)}{\partial x}=400x\left(x^{2}-y\right)+2(x-1)=0\\ \frac{\partial f_{3}(x,y)}{\partial y}=200\left(y-x^{2}\right)=0\end{array},\right.$
at $(x,y)=(1,1)$. So, it has a unique global optimum $(1,1)$. Suppose $\left(x_{0},y_{0}\right)$ is a PNE, then $x_{0}$ and $y_{0}$ are the global minima of differentiable $f_{3}\left(x,y_{0}\right)$ and $f_{3}\left(x_{0},y\right)$, respectively. Then
$\left\{\begin{array}[]{c}\frac{df_{3}\left(x,y_{0}\right)}{dx}=400x^{3}-400xy_{0}+2x-2=0\\ \frac{df_{3}\left(x_{0},y\right)}{dy}=200y-200x_{0}^{2}=0\end{array},\right.$
we have $200x^{3}-200xx_{0}^{2}+x-1=0,y=x_{0}^{2}=y_{0}$. Since $x_{0}$ is one of solutions of equation $0200x^{3}-200xx_{0}^{2}+x-1=$ 0, we conclude that $x_{0}=1,y_{0}=$ 1.
B.4 Function 4
$f_{4}(x,y)=|x-y|-\min(x,y)=\left\{\begin{array}[]{c}x-2y,x>y\\ -x,x=y\\ y-2x,x<y\end{array}\right.$
For any $\left(x_{0},y_{0}\right)$, the function $f_{4}\left(x,y_{0}\right)$ and $f_{4}\left(x_{0},y\right)$ obtain their global minima at $\left(y_{0},y_{0}\right)$ and $\left(x_{0},x_{0}\right)$, respectively. So, the set $\{(x,y)\mid x=y\}$ is the set of PNEs.
Appendix C Convergence Analysis on A Function with Loss of Gradients
Corollary 3: For the Schwefel’s Problem 2.21, $f(\mathbf{x})=\max_{i=1,\ldots,n}\left(\left|x_{i}\right|\right)$, defined on an open set [^19] Here it is implicitly assumed that there is (at least) one global optimum in this open set. $\Omega\subseteq\mathbb{R}^{n}$, the set of global pure Nash equilibria w.r.t. any partition $p=\left\{g_{1},\ldots,g_{m}\right\}$ is $\left\{\mathbf{x},\max_{j\in g_{1}}\left|x_{j}\right|=\cdots=\right.$ $\left.\max_{j\in g_{m}}\left|x_{j}\right|\right\}$. There is a unique strict global Nash equilibrium $\mathbf{x}=(0,\ldots,0)$ w.r.t. any partition set, which equals the global optimum, and vice versa.
Proof: Given $\mathbf{x}=\left(\mathbf{x}_{g_{1}},\cdots,\mathbf{x}_{g_{m}}\right)$ is a PNE w.r.t. any partition $p=\left\{g_{1},\ldots,g_{m}\right\}$ of $f(\mathbf{x})=\max_{i=1,\ldots n}\left(\left|x_{i}\right|\right)$, owing to the definition of PNE, we have $\mathbf{x}_{g_{i}}\in\mathbb{R}^{\left|g_{i}\right|}$ for each $g_{i}\in p,\forall i\in\{1,\ldots,m\}$, satisfies
$f\left(\mathbf{x}_{g_{i}},\mathbf{x}_{\neq g_{i}}\right)\leq f\left(\mathbf{x}_{g_{i}}^{\sim},\mathbf{x}_{\neq g_{i}}\right),\forall\mathbf{x}_{g_{i}}^{\sim}\in\mathbb{R}^{\left|g_{i}\right|}\backslash\left\{\mathbf{x}_{g_{i}}\right\},$
namely, for any $i\in\{1,\cdots,m\},\mathbf{x}_{g_{i}}^{\sim}\in\mathbb{R}^{|g_{i}|}\backslash\left\{\mathbf{x}_{g_{i}}\right\},\max\left\{\max_{j\in g_{i}}\left|x_{j}\right|,\max_{j\in\neq g_{i}}\left|x_{j}\right|\right\}\leq\max\left\{\max_{j\in g_{i}}\left|x_{g_{i}}^{\sim}\right|\right.$, $\left.\max_{j\in\neq g_{i}}\left|x_{j}\right|\right\}$, only if, $\max\left\{\max_{j\in g_{i}}\left|x_{j}\right|,\max_{j\in\neq g_{i}}\left|x_{j}\right|\right\}\leq\min\left\{\max_{j\in g_{i}}\left|x_{g_{i}}^{\sim}\right|,\max_{j\in\neq g_{i}}\left|x_{j}\right|\right\}$, only if,
$\left\{\begin{array}[]{c}\max_{j\in g_{i}}\left|x_{j}\right|\leq\min\left\{\max_{j\in g_{i}}\left|x_{g_{i}}^{\sim}\right|\right\}\\ \max_{j\in g_{i}}\left|x_{j}\right|\leq\max_{j\in\neq g_{i}}\left|x_{j}\right|\\ \max_{j\in\neq g_{i}}\left|x_{j}\right|\leq\min\left\{\max_{j\in g_{i}}\left|x_{g_{i}}^{\sim}\right|\right\}\end{array}\right.$
owing to the definition of PNE,
$\max_{j\in g_{i}}\left|x_{j}\right|\leq\min\left\{\max_{j\in g_{i}}\left|x_{g_{i}}^{\sim}\right|\right\}$
so, we only need for any $i\in\{1,\cdots,m\}$,
$\left\{\begin{array}[]{c}\max_{j\in g_{i}}\left|x_{j}\right|\leq\max_{j\in\neq g_{i}}\left|x_{j}\right|\\ \max_{j\in\neq g_{i}}\left|x_{j}\right|\leq\max_{j\in g_{i}}\left|x_{j}\right|\end{array}\right.$
we have $\max_{j\in g_{1}}\left|x_{j}\right|=\cdots=\max_{j\in g_{m}}\left|x_{j}\right|$.
Since $f(\mathbf{x})\geq 0,\mathbf{x_{0}}=(0,\cdots,0)$ is a PNE, and $f\left(\mathbf{x_{0}}\right)=f(0,\cdots,0)=\mathbf{0},\mathbf{x_{0}}$ is the unique global optimum, thus it is a unique strict global Nash equilibrium, and vice versa.